OpenStudy (anonymous):

convert the decimal to a fraction 0.12

OpenStudy (anonymous):

12/100=3/25

OpenStudy (anonymous):

0.12=12/100 =3/25

OpenStudy (anonymous):

For every digit in the decimal after the decimal point add a zero to the one in the denominator. 0.12 = 0.12/1 and there are two decimal digits. Hence it is 12/100. Now, Both 12 and 100 are divisible by 4 (Remember HCF or GCD?!) so divide them by 4 and voila! .12 = 3/25

OpenStudy (anonymous):

tks